In a football game a kicker attempts a field goal. The ball remains in contact with the kicker's foot for 0.0513 s, during which time it experiences an acceleration of 275 m/s2. The ball is launched at an angle of 51.0° above the ground. Determine the (a) horizontal and (b) vertical components of the launch velocity.
Answer #1

V = u + at = 0 + (275 x 0.0513) = 14.11 m/s

A)

Horizontal component, Vx = 14.11 cos51 = 8.88 m/s

B)

Vertical component, Vy = 14.11 sin51 = 10.97 m/s

Comment in case any doubt please rate my answer ....